how to solve this x2y''-xy'+y=2x by variation of parameter


1
Expert's answer
2021-06-28T16:40:46-0400

This equation is euler-cauchy equation.

Substitute "y=x^{m}" into the homogenous equation.


"y'=mx^{m-1}"

"y''=m(m-1)x^{m-2}"

"x^2(m(m-1)x^{m-2})-x(mx^{m-1})+x^m=0"

"x^m(m^2-m-m+1)=0"

"x^m(m-1)^2=0"

"m=1"

Hence the solution of the homogenous equation is


"y=c_1x+c_2x\ln x"

Use method of variations of parameters "c_1=c_1(x), c_2=c_2(x)" 


"y'=c_1+c_1'x+c_2\ln x+c_2+c_2'x\ln x"

Suppose

"c_1'x+c_2'x\ln x=0"

Then


"y'=c_1+c_2\ln x+c_2"

"y''=c_1'+c_2\dfrac{1}{x}+c_2'\ln x+c_2'"

"x^2c_1'+xc_2+x^2c_2'\ln x+x^2c_2'-xc_1-xc_2\ln x"

"-xc_2+c_1x+c_2x\ln x=2x"

We have the system


"c_1'x+c_2'x\ln x=0"


"x^2c_1'+x^2c_2'\ln x+x^2c_2'=2x"

Then


"c_1'=-c_2'\ln x"

"-x^2c_2'\ln x+x^2c_2'\ln x+x^2c_2'=2x"



"c_1'=-c_2'\ln x"

"c_2'x=2"

"c_2=\int\dfrac{2dx}{x}"

"c_1=-\int\dfrac{2\ln xdx}{x}"

"c_1=-\ln^2(x)+C_3"

"c_2=2\ln(x)+C_4"


"y=-x\ln^2(x)+C_3x+2x\ln^2(x)+C_4x\ln(x)"

"y=x\ln^2(x)+C_3x+C_4x\ln(x)"